OpenStudy (zenmo):

Find the first 3 nonzero terms of Maclaurin series, question on understanding the solution.

OpenStudy (zenmo):

I circled the part that I don't understand with a blue circle, why is a 2 being added there?

zepdrix (zepdrix):

Hmm <.<

zepdrix (zepdrix):

Oh oh, because when you multiply these expansions together, you also get x^4 from other combinations like x^2 * x^2. There is no x^1 * x^3 in these series, so we don't have to worry about anything else. Just the x^4 from the first series, the x^4 from the second series, and the x^2 from each of them multiplied together, \(\large\rm -x^2\cdot(-2x^2)\)

OpenStudy (zenmo):

\[(-x^2)*(-2x^2)=2x^2\] would that be correct?

OpenStudy (zenmo):

I'm probably missing something, I don't venture well in these areas. Haha

zepdrix (zepdrix):

\[\large\rm -x^2\cdot(-2x^2)\quad=\quad 2x^{2+2}\]

OpenStudy (zenmo):

I gotten the same answer by doing the long way by doing \[(1-x^2+\frac{ 1 }{ 2 }x^4)(1-2x^2+\frac{ 2 }{ 3 }x^4)\]

OpenStudy (zenmo):

I still don't see how a 2 is being added for \[(\frac{ 2 }{ 3 }+2+\frac{ 1 }{ 2 })x^4\]

OpenStudy (zenmo):

I can see for ..x^2 by adding the two coefficients \[1-[(-1)+(-2)]x^3... = 1 -3x^3+..\]

OpenStudy (zenmo):

ops by x^3 i mean x^2

zepdrix (zepdrix):

You don't have to expand this all the way out.\[\large\rm \left(1-x^2+\frac12x^4\right)\left(1-2x^2+\frac23x^4\right)\]We only want the x^4 powers that come out of this. So we can cherry pick. \(\large\rm 1\cdot\frac23x^4\) \(\large\rm \frac12x^4\cdot1\) \(\large\rm -x^2\cdot(-2x^2)\) Those are the three multiplications which will give us x^4.

zepdrix (zepdrix):

Are you a little rusty on exponent rule maybe? :o Did you understand that multiplication I did earlier?

OpenStudy (zenmo):

I know how to do exponent rules

zepdrix (zepdrix):

So those three multiplications give us, \(\large\rm \frac23x^4\) \(\large\rm \frac12x^4\) \(\large\rm 2x^4\) See where the 2 is coming from? :o

OpenStudy (zenmo):

Oh, I see it now, by doing the foil method or handing picking like you said, you will get 2x^4 from there

OpenStudy (zenmo):

So when you combine like terms, that is where the 2 comes from

OpenStudy (zenmo):

Thanks for the help! :)